OpenStudy (anonymous):

Please help A cell phone plan consists of a monthly charge of $39.99 with the first 200 minutes free each month. An additional charge of $.05 per min. occurs for each min. over 200. HOW MANY ADDITIONAL MIN. WERE USED OVER THE FIRST 200? What is the main idea? What do you know and what do you need to find?

OpenStudy (anonymous):

I'm assuming you forgot to put what was the actual charge for the month.

OpenStudy (dumbcow):

Let C be total monthly cost, E be extra cost due to additional minutes, T is additional minutes C-39.99 = E E = 0.05*T For this problem you want to solve for T T = E/.05 = (C-39.99)/.05